OpenStudy (anonymous):

A local hamburger shop sold a combined total of 600 hamburgers and cheeseburgers on Friday. There were 50 more cheeseburgers sold than hamburgers. How many hamburgers were sold on Friday?

OpenStudy (across):

Let \(H\) stand for hamburgers and \(C\) for cheeseburgers. The first sentence tells us that\[H+C=600.\]The second sentence tells us that\[C=H+50.\]This is a system of two equations. Do you know how to solve it?

OpenStudy (across):

Solve for \(H\), that is.\[\]

OpenStudy (anonymous):

550 for the first one? not really sure =/

OpenStudy (across):

Let’s walk through it. :) We came up with the following system:\[H+C=600,\]\[C=H+50.\]Since we want to solve for \(H\), let’s first rearrange the first equation a little bit (with the intent of solving for \(H\)):\[H=600-C.\]Are you with me thus far?

OpenStudy (anonymous):

yea

OpenStudy (across):

We now have the following system:\[H=600-C,\]\[C=H+50.\]Since we still want to find \(H\), we want to get rid of \(C\). Effectively enough, the second equation tells us what \(C\) is. We now plug it into the first equation, as follows:\[H=600-[H+50].\]Are you still with me?

OpenStudy (anonymous):

not really =/

OpenStudy (across):

Let me break it down for you into two steps: 1. We ended up with the following system:\[H=600-C,\]\[C=H+50.\]2. We want to solve for \(H\), so we want to get rid of \(C\). The first equation has a \(C\) on the right hand side, and from the second equation, we know that \(C=H+50\). So we basically “plug” the second equation into the first one, like this: \[H=600-C\implies H=600-(H+50)\](since \(C=H+50\)).

OpenStudy (anonymous):

can you please provide the answer pls i'm so lost with this

OpenStudy (across):

Okay… … all we have to do now is simplify the last expression we obtained:\[H=600-(H+50)\]\[\implies H=600-H-50\]\[\implies2H=550\]\[\implies H= 275.\]That’s how many hamburgers were sold.

OpenStudy (anonymous):

Alright thanks so much!